OpenStudy (anonymous):

In the US about 15% of the population is 65 years or older. It is estimated that 90% of seniors get a flu shot. It is also estimated that 45% of non-seniors get the flu shot. If a US citizen is selected at random: A) what is the probability that they are under 65years old? B)what is the probability they are over 65 and did not get a flu shot? C) what is the probability that they got a flu shot?

OpenStudy (kropot72):

A)\[P(age \ge65)=0.15\] \[P(age <65)=1.00-0.15=you\ can\ calculate\]

OpenStudy (kropot72):

B) I assume that the question should be: "what is the probability they are 65 or older and did not get a flu shot?" Let A be the event '65 years or older' and let B be the event 'senior did not get a flu shot'. Then: P(A) = 0.15 P(B) = 1.0 - 0.9 = 0.1 \[P(A \cap B)=P(A) \times P(B)=0.15\times0.1=you\ can\ calculate\]

OpenStudy (kropot72):

C) Let C be the event 'senior got a flu shot'. \[P(C)=0.15\times0.9\] Let D be the event 'non-senior got a flu shot'. \[P(D)=0.85\times0.45\] Let E be the event 'US citizen got a flu shot'. Events C and D are mutually exclusive, therefore: \[P(E)=P(C)+P(D)=(0.15\times0.9)+(0.85\times0.45)=you\ can\ calculate\]

OpenStudy (anonymous):

Thank you! and yes I can do the simple calcualtions or my calculator can :-)

OpenStudy (kropot72):

Cool! :)